# Game-based Platforms for Artificial Intelligence Research

### v3

@@ IV-A Game Tuning
Game tuning is vital since it directly affects players’ experience. Unsuitable game parameter settings may lead to game unbalance or low quality, where players find it hard to play the game or feel bored. Generally, all the games introduced in Section III that involve configurable parameters can be tuned with search-based methods for different purposes, such as adjusting the game difficulty. For instance, FightingICE [56] has been used to study the dynamic difficulty and battle sound adjustment [115]. VGC [35] allows adjusting the attributes of Pokémons, called meta-game balance. Compared to other platforms, it specifically facilitates the investigation of enabling fair competition in gameplay. AIGD2 [55], an open-source game similar to Planet Wars, allows to explore the parameter space, which can significantly affect the gameplay via GVGAI agents [116].
IV-B Content Generation
@@
Content generation is involved with applying AI techniques to create new game content. The generated content types include 2D level, 3D level, narrative, rhythm chart and card, following Liu et al. [15]. Different content types are represented in different data structures, raising various challenges to content generation research.
IV-B1 2D Map & Level
@@
Game level is the most commonly generated content type. A 2D level is usually presented in a plane, where agents move in two dimensions. Mario-AI-Framework [50], a fully-Java copy of Super Mario Bros., provides a tile-based representation of game levels. Fig. 5 shows examples of Mario levels generated by [117], tested and visualised in Mario-AI-Framework [50]. Hierarchical Open Graph 2 (HOG2) [94] includes several grid-based puzzle games such as Snakebird [118] for exhaustive PCG [119]. Morai-Maker [98] is a Unity3D-based platform for Super Mario Bros., which provides three built-in machine learning-based 2D level generators. Co-creation between humans and AI is supported by the platform with a visible game level editor [98], which distinguishes Morai-Maker from other Super Mario Bros. platforms. Procedural Zelda [95] is a grid-based game platform for level, combat and puzzle generation. Both offline and online control are supported during the generation with data logging capabilities. PCGRL gym [99] implements RL environments for grid-based level generation with the Gym interface. Binary, Zelda and Sokoban problems are integrated into PCGRL gym. micro-rct [100] is a theme park management simulator, inspired by RollerCoaster Tycoon 2. In [100], researchers use a controllable PCGRL to generate game map. micro-rct is used to evaluate the controllability of PCGRL. Unlike other platforms which serve for search-based, supervised learning-based, or unsupervised learning-based content generation algorithms, PCGRL gym [100] provides a standardised benchmark for RL-based PCG, however, no accessible playing interface for human players is provided. Science Birds [97] provides a generative platform, which can load XML-based level files to play Angry Birds, a physics-based puzzle video game, which requires the level components’ physical feasibility and stability. Generating levels for physical-based puzzle games with AI is complex since more physical constraints, larger state space and action space should be considered. Angry Birds AI Level Generation Competition [52] uses Science Birds as its simulator.
@@
Recently, LLMs have obtained much attention [120, 121]. Chatgpt4PCG [122] competition aims at using ChatGPT to generate game levels of Angry Birds. Although no specific environment API is provided, an example prompt is introduced. Chatgpt4PCG enables level generation via natural language, which is more user-friendly for beginners. Meanwhile, the investigation into generating levels with ChatGPT can help researchers to better understand the ability and limitations of LLMs. Talakat [96] is a bullet hell game engine which defines a domain-specific language for bullet hell levels, supporting the grammar evolution of bullet hell game content. Keiki [101] is another bullet hell type game engine which provides a native Python interface describing bullet hell barrage. While Talakat represents barrages as scripts, Keiki provides a function to convert barrage scripts into vector-based representations, making it convenient for ML-based content generators.
IV-B2 3D Map & Level
@@ IV-B3 Narrative
IV-B4 Rhythm Chart
Rhythm game is a special 2D game that aims to hit charts at the right time. To optimise player experience, the generated charts need to be synchronous with the rhythm of the music. StepMania [106] is a free dance and rhythm game engine based on game Dance Dance Revolution that supports customised rhythm charts. StepMania has been researched in level generation i.e. chart generation [126], which is related to the music information retrieval field using long short-term memory (LSTM) and C-LSTM model. StepMania can be used in difficulty estimation [127] as well. A beatmap and rhythm chart can be generated automatically according to the music [128, 129] by a rhythm game platform OSU! [107].
IV-B5 Board & Card
@@
IV-C Other Resources for AI-involved Game Design
@@
IV-D Guidance for Designing Games
We summarise the games and platforms for game design in Tab. III. A majority of game design platforms focus on content generation. Programming languages vary across the games and platforms, with half of them supporting Java. Considering the trend of an increasing portion of using Python in AI research, it is recommended to use those platforms with cross-language programming. For example, researchers can access Java programs in Python via Jpype [^3] https://github.com/jpype-project/jpype/.
@@ IV-D Guidance for Designing Games
For other types of content, there are very limited numbers of platforms to use. Researchers can directly choose the corresponding platform according to the interested content type with the help of Tab. III. FightingICE, AIGD2 and VGC are platforms that can be used for studying game tuning and balancing, but with different focus on fighting games, RTS games and Pokémon, respectively.
V Discussion and Outlook
@@
V-A Challenges from Games
@@
V-A1 Curse of High Dimensions
@@
V-A2 Poor Exploration
AI techniques such as RL usually need a large amount of interaction data to train policies. It is hard to directly access such a “dataset”. Instead, we need to collect the data, denoted as experience, during training, which leads to a tradeoff between exploration and exploitation. The tradeoff happens with a limited computational budget since we hardly go through all possible states and actions. Montezuma’s revenge epitomises the games where AI techniques suffer from the tradeoff [136], which brings a deep thought about the exploration strategy. Integrating the exploration signal into the reward function has been shown efficient to tackle the challenge, such as count-based exploration [137], intrinsic curiosity [138] and novelty search [139]. While the above reward shaping may lead to the optimal policy variance, population-based exploration [140] and noisy exploration [136, 141] that perturb the policy parameter space present their advantage for the diverse exploration.
V-A3 Partial Observability
@@
V-A4 Generalisation
Although AI techniques such as deep RL have achieved a competitive level beyond humans in games including Go [5] and Atari [68], their generalisation ability is questionable since they are usually only trained on a given game or level. The generalisation considers if the agent can still play well in some unseen games and levels without overfitting the training environment [143]. Many game-based environments are involved with the topic. GVGAI [71, 72] focus on the performance of agents when trained on some game levels but tested on unseen game levels. The unseen test game levels have different game maps or layouts, but their mechanism is the same [77]. Both planning and learning algorithms are supported by the GVGAI environment [71, 72]. Procgen [77] also provides diverse game scenarios, which test the AI agents by randomly changing the initial position of the avatars.
V-A5 Multi-agent Games
@@ V-B1 Game Testing by AI
Moreover, games often require frequent updates to introduce new content and address gameplay issues. This necessity can be both a source of excitement and frustration for players. The critical challenge lies in testing updated content and maintaining game balance. Such studied [152, 153] have been conducted on multiple games such as VGC [35] and StarCraft II [6]. Advanced testing methodologies are capable of adapting to the rapidly changing landscape of game development, which enhances the player experience without compromising game quality or balance.
V-B2 Game Design by AI
In games, levels is usually the most fantastic part, which represents the effort and inspiration of game designers. However, creating levels requires a significant investment of time and human resources. AI has the potential of accelerating this process and can even serve as a substitute for human designers [12, 13, 14, 15, 16]. Besides, AI can be applied to generate music and images in games. For example, Ferreira et al [154]. proposed a DL-based music composition system to generate music for tabletop role-playing games. Hong et al. applied multi-discriminator generative adversarial networks to generate game sprites [155]. On the other hand, other 2D and 3D content generation methods [156], not limited to AI approaches, can also be tested and applied to games. Moreover, game stories and programming codes can be generated via AI technologies. For example, Ubisoft has recently proposed Ghostwriter to assist game designers to write game stories and dialogues [157]. Ghostwriter receives feedbacks from designers to construct conversations between NPCs, instead of replacing designers directly.
@@ V-C Games for AI
V-C1 Open-world Games for Artificial General Intelligence
General (video) game playing has been extensively studied using multi-game platforms [62, 71, 72, 92]. However, the objectives or tasks in such games are often well-defined and known. Open-world games, e.g., MineRL [59], pose new challenges for AI to autonomously explore the Minecraft world without specific aims, including the requirement of accomplishing a series or set of tasks instead of one or two, with inefficient data sampling and indeterminate reward function [59]. The nature of open-world games provides an ideal platform to develop and evaluate artificial general intelligence. They can be considered as sandboxes, where AI agents are not only tested on some specific tasks such as navigating and building, but also are required to interact with the dynamic environment. The diverse tasks such as farming, hunting and crafting in Minecraft enable AI to live as humans. This case is pushing the boundaries of AI towards achieving artificial general intelligence, as they mimic the complexity and unexpected of real-world tasks and decision-making.
V-C2 Games as Reflection of Real-world Scenarios
@@
V-C3 Emerging Environments for AI Creativity
@@
V-D Under-explored Game Genres
@@
Although there are many available games, they share a large number of similar mechanisms with only differing in sprite styles. Many platforms for AI research are based on classic games like Atari [68]. Although the simplified representation diminishes some engineering problems, this may not lead to a more delicate game mechanism thereby promoting AI research. Only a few modern commercial games such as StarCraft II [6], and Dota 2 [7], provide accessible APIs or lightweight simulators.
@@
Extensive effort has to be put into solving engineering problems such as controlling avatars, which is a critical issue in conducting AI research on commercial games. A possible solution for this is capturing the screening as input. AI agent directly accesses the screen frames and plays the game via a virtual keyboard and mouse. As new games are developed, challenges arise for AI research. It is expected that more commercial games provide accessible APIs that AI can interact with. Similar concerns exist in game-based platforms as most platforms for content generation are based on rather classic video games like Super Mario Bros. [50], Angry Birds [97, 52] and Minecraft [59, 60]. Though using classical games makes it convenient for research, it can prevent AI-involved game design from being used in practice.
V-E Lower Barriers for Building New Games
@@
There are some professional game engines to build simulators for machine learning such as Unity [80], Unreal Engine [^5] https://github.com/zfw1226/gym-unrealcv and Godot [160]. Developing a domain-specific platform for a game is not easy. Some simulators can only be used in specific scenarios, and if the AI algorithm needs to be used in other situations, a new simulator may need to be built, which is time-consuming. One way to make simulators more versatile is by applying description languages. VGDL [91] is a good example. Games written by VGDL can be compiled by a specific compiler such as GVGAI [71, 72, 92] and UnityVGDL [161]. Another way is using general platforms, especially professional game engines to build simulators such as Unity [80], Unreal Engine and Godot [160]. Using professional game engines, simulators can be built much easier and run faster. With remote procedure call, algorithms do not need to be implemented with the same programming language as simulators do, which makes it possible to use general platforms for AI research, reducing engineering issues.
V-F Large Language Models for Games
@@
Recently, the emergence of LLMs such as ChatGPT [145] narrows the gap of game generation [121]. Using natural languages, i.e., interacting with LLMs, everyone can create their own game. Some other applications of LLMs for assisting game development such as level generation and dialogue generation also thrive. MarioGPT [20, 21] provides an end-to-end Mario level generator through a fine-tuned GPT2. Users can write some simple prompts such as “many pipes, some enemies” to generate game levels with their preference. The work of [19] validates the effectiveness of LLMs for generating Sokoban levels from the perspective of playability, novelty, diversity and accuracy.
@@
Instead of assisting game content generation, LLM itself can also be regarded as a part of games. NPCs have been playing an essential role in games. Traditionally, NPCs have been confined to fixed scripts, limiting them to repetitive conversations that do not interact with or affect players’ strategies. LLMs, however, provide the potential of human-like NPCs, who may act and respond according to the behaviours of players. For example, Park et al. [17] adopt ChatGPT for the generative agent design. Players directly communicate with NPCs controlled by ChatGPT. Kumaran et al. [18] propose SceneCraft based on LLM for narrative generation in a 3D game. In addition to dialogue, SceneCraft generates emotion and gestures of NPCs during the interactions with human players. The work of [162] shows the possibility of developing software including games using multiple LLM agents.
@@
Notably, LLMs also show the potential of playing games. Wu et al. [22] use GPT-4 to read a paper about game Spring, and then directly apply GPT-4 to play the game. According to [22], GPT-4 presents outstanding performance than advanced RL algorithms without training. It is widely believed that LLMs, with their remarkable capabilities and capacity for generalisation, are driving the future direction of game and game AI development.
V-G Roles of Games in Interdisciplinary Research in AI

### v4

@@ 4.1 Game Tuning
Game tuning is vital since it directly affects players’ experience. Unsuitable game parameter settings may lead to game unbalance or low quality, where players find it hard to play the game or feel bored. Generally, all the games introduced in Section 3 that involve configurable parameters can be tuned with search-based methods for different purposes, such as adjusting the game difficulty. For instance, FightingICE [77] has been used to study the dynamic difficulty and battle sound adjustment [151]. AIGD2 [75], an open-source game similar to Planet Wars, allows to explore the parameter space, which can significantly affect the gameplay via GVGAI agents [152]. VGC [54] allows adjusting the attributes of Pokémons, called meta-game balance. Compared to other platforms, it specifically facilitates the investigation of enabling fair competition in gameplay. RaidEnv [129] aims to optimise boss raid scenarios. Skills and statistics like health and attack can be adjusted by AI techniques, meeting the needs of different situations.
4.2 Content Generation
@@
Content generation is involved with applying AI techniques to create new game content. The generated content types include 2D levels, 3D levels, narratives, rhythm charts and cards, following Liu et al. [6]. Different content types are represented in different data structures, raising various challenges to content generation research.
4.2.1 2D Map & Level
Game level is the most commonly generated content type. A 2D level is usually presented in a plane, where agents move in two dimensions. Science Birds [130] provides a generative platform, which can load XML-based level files to play Angry Birds, a physics-based puzzle video game, which requires the level components’ physical feasibility and stability. Generating levels for physical-based puzzle games with AI is complex since more physical constraints, larger state space and action space should be considered. Angry Birds AI Level Generation Competition [78] uses Science Birds as its simulator.
@@
Mario-AI-Framework [72], a fully-Java copy of Super Mario Bros., provides a tile-based representation of game levels. Fig. 5 shows examples of Mario levels generated by [153], tested and visualised in Mario-AI-Framework [72]. Morai-Maker [131] is a Unity3D-based platform for Super Mario Bros., which provides three built-in ML-based 2D level generators. Co-creation between humans and AI is supported by the platform with a visible game level editor [131], which distinguishes Morai-Maker from other Super Mario Bros. platforms. Hierarchical Open Graph 2 (HOG2) [132] includes several grid-based puzzle games such as Snakebird [154] for exhaustive PCG [155]. Procedural Zelda [133] is a grid-based game platform for level, combat and puzzle generation. Both offline and online control are supported during the generation with data logging capabilities.
Talakat [134] is a bullet hell game engine which defines a domain-specific language for bullet hell levels, supporting the grammar evolution of bullet hell game content. Keiki [135] is another bullet hell type game engine which provides a native Python interface describing bullet hell barrage. While Talakat represents barrages as scripts, Keiki provides a function to convert barrage scripts into vector-based representations, making it convenient for ML-based content generators.
Procedural content generation via reinforcement learning (PCGRL) gym [136] implements RL environments for grid-based level generation with the Gym interface. Binary, Zelda and Sokoban are integrated into PCGRL gym. micro-rct [137] is a theme park management simulator, inspired by RollerCoaster Tycoon 2. In Earle et al. [137], researchers use a controllable PCGRL to generate game maps. micro-rct is used to evaluate the controllability of PCGRL. Unlike other platforms which serve for search-based, supervised learning-based, or unsupervised learning-based content generation algorithms, PCGRL gym [137] provides a standardised benchmark for RL-based PCG, however, no accessible playing interface for human players is provided.
@@
Recently, LLMs have obtained much attention [156, 157]. Chatgpt4PCG [158] competition aims at using ChatGPT to generate game levels of Angry Birds. Although no specific environment API is provided, an example prompt is introduced. Chatgpt4PCG enables level generation via natural language, which is more user-friendly for beginners. Meanwhile, the investigation into generating levels with ChatGPT can help researchers to better understand the ability and limitations of LLMs.
4.2.2 3D Map & Level
@@ 4.2.3 Narrative
4.2.4 Rhythm Chart
Rhythm game is a special 2D game that aims to hit charts at the right time. To optimise player experience, the generated charts need to be synchronous with the rhythm of the music. StepMania [142] is a free dance and rhythm game engine based on game Dance Dance Revolution that supports customised rhythm charts. StepMania has been researched in level generation, i.e., chart generation [162], which is related to the music information retrieval field using long short-term memory (LSTM) and C-LSTM model. StepMania can be used in difficulty estimation [163] as well. A beatmap and rhythm chart can be generated automatically according to the music [164, 165] by a rhythm game platform OSU! [143].
4.2.5 Board & Card
@@
4.3 Other Resources for AI-involved Game Design
@@
4.4 Guidance for Designing Games
Tab. 3 summarises the games and platforms for game design, in which how AI can be utilised can be checked in column Type. Programming languages vary across the games and platforms, with half of them supporting Java. Considering the trend of an increasing portion of using Python in AI research, it is recommended to use those platforms with cross-language programming. For example, researchers can access Java programs in Python via Jpype [^3] https://github.com/jpype-project/jpype/.
Some specific games, such as FightICE and VGC, can be used for studying game balance and skill discovery via parameter tuning techniques [170, 171, 172]. A majority of games for design focus on procedurally generating content such as 2D and 3D levels. AI techniques including rule-based, search-based, ML-based and DL-based methods have been widely used for PCG using games listed in Tab. 3 [22, 23, 5, 24, 25, 6, 26, 27]. For example, bullet hell and maze level can be realised by grammar evolution [134, 173] and quality diversity search [174]. Real-time generation of 2D levels of Super Mario Bros. represented by latent vectors also benefits from RL [175, 153]. These games also serve as benchmarks for generative models like generative adversarial networks [176], diffusion models [177] and LLMs [32]. In addition, supervised learning approaches have been applied to rhythm chart generation [162, 165] using StepMania and OSU!, as well as card generation [144] based on XMage. Notably, games for content generation like Procedural Zelda are also involved with adjusting game difficulty [178], and tuning parameters of games or agents [152, 179, 180].
@@ 4.4 Guidance for Designing Games
For other types of content, there are very limited numbers of platforms to use. Researchers can directly choose the corresponding platform according to the interested content type with the help of Tab. 3. FightingICE, AIGD2 and VGC are platforms that can be used for studying game tuning and balancing, but with different focus on fighting games, RTS games and Pokémon, respectively.
5 Discussion and Outlook
@@
5.1 Challenges from Games
@@
5.1.1 Curse of High Dimensions
@@
5.1.2 Poor Exploration
AI techniques such as RL usually need a large amount of interaction data to train policies. It is hard to directly access such a “dataset”. Instead, we need to collect the data, denoted as experience, during training, which leads to a tradeoff between exploration and exploitation. The tradeoff happens with a limited computational budget since we hardly go through all possible states and actions. Montezuma’s revenge epitomises the games where AI techniques suffer from the tradeoff [183], which brings a deep thought about the exploration strategy. Integrating the exploration signal into the reward function has been shown efficient to tackle the challenge, such as count-based exploration [122], intrinsic curiosity [184] and novelty search [185]. While the above reward shaping may lead to the optimal policy variance, population-based exploration [186] and noisy exploration [183, 187] that perturb the policy parameter space, present their advantage for the diverse exploration.
5.1.3 Partial Observability
@@
5.1.4 Generalisation
Although AI techniques such as deep RL have achieved a competitive level beyond humans in games including Go [14] and Atari [93], their generalisation ability is questionable since they are usually only trained on a single game or level. The generalisation considers if the agent can still play well in some unseen games and levels without overfitting the training environment [189]. Many game-based environments are involved with the topic. GVGAI [95, 96] focus on the performance of agents when trained on some game levels but tested on unseen game levels. The unseen test game levels have different game maps or layouts, but their mechanism is the same [102]. Both planning and learning algorithms are supported by the GVGAI environment [95, 96]. Procgen [102] also provides diverse game scenarios, which test the AI agents by randomly changing the initial position of the avatars.
5.1.5 Multi-agent Games
@@ 5.2.1 Game Testing by AI
Moreover, games often require frequent updates to introduce new content and address gameplay issues. This necessity can be both a source of excitement and frustration for players. The critical challenge lies in testing updated content and maintaining game balance. Such studies [198, 170] have been conducted on multiple games such as VGC [54] and StarCraft II [15]. Advanced testing methodologies are capable of adapting to the rapidly changing landscape of game development, which enhances the player experience without compromising game quality or balance.
5.2.2 Game Design by AI
In games, levels is usually the most fantastic part, which represents the effort and inspiration of game designers. However, creating levels requires a significant investment of time and human resources. AI has the potential of accelerating this process and can even serve as a substitute for human designers [22, 23, 25, 6, 26].
Besides, AI can be applied to generate music and images in games. For example, Ferreira et al [199]. proposed a DL-based music composition system to generate music for tabletop role-playing games. Hong et al. applied multi-discriminator generative adversarial networks to generate game sprites [200]. On the other hand, other 2D and 3D content generation methods [201], not limited to AI approaches, can also be tested and applied to games. Moreover, game stories and programming codes can be generated via AI technologies. For example, Ubisoft has recently proposed Ghostwriter to assist game designers to write game stories and dialogues [202]. Ghostwriter receives feedbacks from designers to construct conversations between NPCs, instead of replacing designers directly.
@@ 5.3 Games for AI
5.3.1 Open-world Games for Artificial General Intelligence
General (video) game playing has been extensively studied using multi-game platforms [87, 95, 96, 127]. However, the objectives or tasks in such games are often well-defined and known. Open-world games, e.g., MineRL [82], pose new challenges for AI to autonomously explore the Minecraft world without specific aims, including the requirement of accomplishing a series or set of tasks instead of one or two, with inefficient data sampling and indeterminate reward function [82]. The nature of open-world games provides an ideal platform to develop and evaluate artificial general intelligence. They can be considered as sandboxes, where AI agents are not only tested on some specific tasks such as navigating and building, but also are required to interact with the dynamic environment. The diverse tasks such as farming, hunting and crafting in Minecraft enable AI to live as humans. This case is pushing the boundaries of AI towards achieving artificial general intelligence, as it mimics the complexity and unexpected cases of real-world tasks and decision-making.
5.3.2 Games as Reflection of Real-world Scenarios
@@
5.3.3 Emerging Environments for AI Creativity
@@
5.4 Under-explored Game Genres
@@
Although there are many available games, they share a large number of similar mechanisms with only differences in sprite styles. Many platforms for AI research are based on classic games like Atari [93]. Although the simplified representation diminishes some engineering problems, it is hard to construct more delicate game mechanisms. Only a few modern commercial games such as StarCraft II [15], and Dota 2 [16], provide accessible APIs or lightweight simulators.
@@
Extensive effort has to be put into solving engineering problems such as controlling avatars, which is a critical issue in conducting AI research on commercial games. A possible solution for this is capturing the screening as input. AI agent directly accesses the screen frames and plays the game via a virtual keyboard and mouse. As new games are developed, challenges arise for AI research. It is expected that more commercial games provide accessible APIs that AI can interact with. Similar concerns exist in game-based platforms as most platforms for content generation are based on rather classic video games like Super Mario Bros. [72], Angry Birds [130, 78] and Minecraft [82, 83]. Though using classical games makes it convenient for research, it prevents AI-involved game design from being used in practice.
5.5 Lower Barriers for Building New Games
@@
There are some professional game engines to build simulators for ML research such as Unity [106], Unreal Engine [^5] https://github.com/zfw1226/gym-unrealcv and Godot [205]. Developing a domain-specific platform for a game is not easy. Some simulators can only be used in specific scenarios, and if the AI algorithm needs to be used in other situations, building a new simulator may be necessary, which is time-consuming. One way to make simulators more versatile is by applying description languages. VGDL [117] is a good example. Games written by VGDL can be compiled by a specific compiler such as GVGAI [95, 96, 127] and UnityVGDL [206]. Another way is using general platforms, especially professional game engines to build simulators such as Unity [106], Unreal Engine and Godot [205]. Using professional game engines, simulators can be built much easier and run faster. With remote procedure call, algorithms do not need to be implemented with the same programming language as simulators do, which makes it possible to use general platforms for AI research, reducing engineering issues.
5.6 Large Language Models for Games
@@
Recently, the emergence of LLMs such as ChatGPT [191] narrows the gap of game generation [157]. Using natural languages, i.e., interacting with LLMs, everyone can create their own games. Some other applications of LLMs for assisting game development such as level generation and dialogue generation also thrive. MarioGPT [31, 32] provides an end-to-end Mario level generator through a fine-tuned GPT2. Users can write some simple prompts such as “many pipes, some enemies” to generate game levels with their preference. The work of [30] validates the effectiveness of LLMs for generating Sokoban levels from the perspective of playability, novelty, diversity and accuracy.
@@
Instead of assisting game content generation, LLMs themselves can also be regarded as a part of games. NPCs have been playing an essential role in games. Traditionally, NPCs have been confined to fixed scripts, limiting them to repetitive conversations that do not interact with or affect players’ strategies. LLMs, however, provide the potential of human-like NPCs, who may act and respond according to the behaviours of players. For example, Park et al. [28] adopt ChatGPT for the generative agent design. Players directly communicate with NPCs controlled by ChatGPT. Kumaran et al. [29] propose SceneCraft based on LLMs for narrative generation in a 3D game. In addition to dialogue, SceneCraft [29] generates emotion and gestures of NPCs during the interactions with human players. The work of [207] shows the possibility of developing software including games using multiple LLM agents.
@@
Notably, LLMs also show the potential of playing games. Wu et al. [33] use GPT-4 to read a paper about game Spring, and then directly apply GPT-4 to play the game. According to [33], GPT-4 presents an outstanding performance than advanced RL algorithms without training. It is widely believed that LLMs, with their remarkable capabilities and capacity for generalisation, are driving the future direction of games and game AI development.
5.7 Roles of Games in Interdisciplinary Research in AI